solve.
6x^4-5x^2+1=0
Please help.

Answer by stanbon(75887)

6x^4-5x^2+1=0
--------------------
Factor:
(6x^2 + 1)(x^2-1) = 0
Solve each factor separately:
6x^2 = -1
x^2 = -1/6
x = +-(1/6)sqrt(6)i
==========================
x^2-1 = 0
(x+1)(x-1) = 0
x = -1 or x= 1
==============================
Cheers,
Stan H.